# Large Language Model Confidence Estimation via Black-Box Access

### v2

@@ 4.1 Confidence Estimation
We use two metrics to evaluate effectiveness of the models: Area under the receiver operating characteristic (AUROC) curve which computes the model’s discrimination ability for various thresholds. The curve is plotted by varying the thresholds of the prediction probabilities of the model and the false positive rate and the true positive rate form the X and the Y axes. The area under this curve is called the AUROC. An accuracy rejection curve can also be plotted by increasing the rejection threshold gradually and plotting the model’s average accuracy at that threshold. The area under this curve is called AUARC [22].
In Table 2, we see that our method quite consistently outperforms all baselines on AUROC. For estimating the confidence of Llama’s responses on TriviaQA, our model is better than the best baseline by $11$ percentage points. We are also able to estimate the confidence on the SQuAD dataset using Mistral by $14$ percentage points better than the closest competitor. Qualitatively similar results are seen for the SQuAD dataset using Flan-ul2 (better by $10$ percentage points) and for the CoQA and NQ datasets using Llama (better by $15$ and $10$ percentage points respectively). Our results on the summarization datasets using LLMs that excel at summarization (viz. Pegasus and BART) we see again that we are either better or at least competitive.
@@
Dataset Source LLM AUROC Self Target LLM 1 AUROC Target LLM 2 AUROC
TriviaQA Llama 0.88 0.94 (Flan-ul2) 0.80 (Mistral)
Flan-ul2 0.94 0.87 (Llama) 0.80 (Mistral)
Mistral 0.81 0.84 (Llama) 0.91 (Flan-ul2)
SQuAD Llama 0.83 0.81 (Flan-ul2) 0.80 (Mistral)
Flan-ul2 0.8 0.79 (Llama) 0.78 (Mistral)
Mistral 0.84 0.82 (Llama) 0.83 (Flan-ul2)
CoQA Llama 0.92 0.79 (Flan-ul2) 0.78 (Mistral)
Flan-ul2 0.87 0.87 (Llama) 0.81 (Mistral)
Mistral 0.81 0.88 (Llama) 0.86 (Flan-ul2)
NQ Llama 0.85 0.91 (Flan-ul2) 0.83 (Mistral)
Flan-ul2 0.93 0.83 (Llama) 0.82 (Mistral)
Mistral 0.83 0.85 (Llama) 0.90 (Flan-ul2)
CNN Pegasus 0.77 0.57 (BART) -
BART 0.57 0.77 (Pegasus) -
XSUM Pegasus 0.73 0.58 (BART) -
BART 0.57 0.71 (Pegasus) -
@@
Dataset Source LLM AUARC Self Target LLM 1 AUARC Target LLM 2 AUARC
TriviaQA Llama 0.83 0.74 (Flan-ul2) 0.64 (Mistral)
Flan-ul2 0.74 0.83 (Llama) 0.64 (Mistral)
Mistral 0.64 0.83 (Llama) 0.73 (Flan-ul2)
SQuAD Llama 0.68 0.62 (Flan-ul2) 0.63 (Mistral)
Flan-ul2 0.96 0.89 (Llama) 0.91 (Mistral)
Mistral 0.96 0.90 (Llama) 0.91 (Flan-ul2)
CoQA Llama 0.71 0.79 (Flan-ul2) 0.61 (Mistral)
Flan-ul2 0.80 0.70 (Llama) 0.61 (Mistral)
Mistral 0.61 0.69 (Llama) 0.79 (Flan-ul2)
NQ Llama 0.45 0.46 (Flan-ul2) 0.42 (Mistral)
Flan-ul2 0.47 0.45 (Llama) 0.42 (Mistral)
Mistral 0.42 0.45 (Llama) 0.46 (Flan-ul2)
CNN Pegasus 0.74 0.34 (BART) -
BART 0.34 0.74 (Pegasus) -
XSUM Pegasus 0.27 0.34 (BART) -
BART 0.35 0.25 (Pegasus) -
@@ 4.2 Confidence Model Interpretability and Transferability
Transferability: Given the commonality between the important features across LLMs for a dataset we now try to test how well does our logistic confidence model for one LLM perform in estimating confidences of another LLM. As seen in Tables 5 and 6 our confidence models are actually quite transferable as they perform comparably or even sometimes better on the other LLMs than the LLM they were built for. This is particularly true for Mistral where, its confidence model performs better for the other two LLMs than itself even coming close in performance to their own confidence models in many cases.
This suggests that we could apply our approach to one LLM and then use the same confidence model to evaluate responses of other LLMs without having to build individual models for them. It would be interesting to further stress test this hypothesis in the future with more LLMs and datasets. Nonetheless, even in the current setup – of five LLMs and six datasets – this observation is interesting and useful.
@@ 5 Discussion
In summary, we have provided an extensible framework for black-box confidence estimation of LLM responses by proposing novel features that are indicative of response correctness. By building an interpretable logistic regression model based on these features we were able to obtain state-of-the-art performance in estimating confidence on six benchmark datasets (CoQA, SQuAD, NQ, TriviaQA, CNN Daily and XSUM) and using five powerful open source LLMs (Llama-2-13b-chat, Mistral-7B-Instruct-v0.2, Flan-ul2, Pegasus-large and BART-large). The interpretability of our confidence model aided in identifying features (viz. SD, SP, EFA,PP) that were instrumental in driving its performance for different LLM-dataset combinations. This led to the interesting realization that many of the features crucial for performance were shared across the confidence models of different LLMs for a dataset. We thus tested if the confidence models generalized across LLMs for a dataset and found that it indeed was the case leading to the interesting possibility of having an universal confidence model trained on just a single LLMs responses, but applied across many others.
@@
Owing to the supervised nature of training the confidence model, one limitation of our approach is that at least some of the model’s generations must be close to the ground truth for us to obtain a reasonable confidence estimator. Another limitation is that the results and insights were obtained based on datasets in English, but these insights might vary when looking at datasets in other languages. More varied tasks and models could be tested upon in the future. We used rouge to test accuracy of generations consistent with previous works, however, rouge, like also other NLP metrics, can be error prone. In terms of broader impact, our approach can be widely applied as it is simple and works with just black-box access to the LLM. Access to logits or internals of the model are not required. However, our estimates although accurate can be imperfect and this should be taken into account when using our approach in high stakes applications involving LLMs. One should also be cognizant of adversaries aware of our features trying to induce misplaced trust in LLMs they create or prefer.
@@
Given the extensibility of our framework, in the future, it would be interesting to add more features as LLMs evolve. One class of such features might be those where the correctness of a response is checked through creating questions that are (causally) related to the original question and context, and seeing how the response varies by asking this question by itself as opposed to in conjunction with the original question and response. Such and other strategies may help in generalizing these confidence estimators also across datasets something that has been seen when we have additional access to logits of LLMs.
@@ Appendix A Prompt Design
  • Flan-ul2 model: Answer the following question in less than 5 words
@@
  • Flan-ul2 model and Llama-2-13b-chat model: Provide an answer in less than 5 words for the following question based on the context below: context: {context} Question: {question} Answer:
@@
  • Flan-ul2 model and Llama-2-13b-chat model: Provide an answer for the following question based on the context below, in less than 5 words:
@@
Prompts for NQ: For all the 3 models we used the following prompt:
@@
Table 7: Percentage of prompt perturbations entailed by the original prompt for the SQuAD dataset. This dataset also has context unlike some of the other Q&A datasets and hence, is a more challenging case of our features to maintain semantics. As can be seen our perturbations produce the intended effect of maintaining the semantics of the original prompt in most cases.
$99.81$% $99.23$% $99.66$% $99.12$%
Table 8: Below we see how the AUROC, AUARC values vary with different number of samples used to train our logistic regression model for some of our datasets. As can be seen our uncertainty estimation procedure is performant even with fewer samples for training.
Dataset LLM $250$ samples $500$ samples $1000$ samples (results in main paper)
TriviaQA Llama 0.83, 0.80 0.86, 0.81 0.88, 0.83
Flan-ul2 0.95, 0.73 0.95, 0.74 0.95, 0.74
Mistral 0.80, 0.63 0.80, 0.63 0.81, 0.64
SQuAD Llama 0.8, 0.65 0.81, 0.66 0.83, 0.68
Flan-ul2 0.76, 0.91 0.78, 0.94 0.8, 0.96
Mistral 0.79, 0.90 0.81, 0.93 0.84, 0.96
CoQA Llama 0.91, 0.70 0.92, 0.71 0.92, 0.71
Flan-ul2 0.86, 0.79 0.87, 0.80 0.87, 0.80
Mistral 0.80, 0.60 0.81, 0.61 0.81, 0.61

### v3

@@ 4.1 Confidence Estimation
We use three metrics to evaluate effectiveness of the models: i) Area under the receiver operating characteristic (AUROC) curve which computes the model’s discrimination ability for various thresholds. The curve is plotted by varying the thresholds of the prediction probabilities of the model and the false positive rate and the true positive rate form the X and the Y axes. The area under this curve is called the AUROC. ii) An accuracy rejection curve can also be plotted by increasing the rejection threshold gradually and plotting the model’s average accuracy at that threshold. The area under this curve is called AUARC [24]. iii) Expected calibration error (ECE) is also reported in Table 5, which measures the discrepancy between accuracy and confidences.
In Table 3, we see that our method quite consistently outperforms all baselines on AUROC. This is also seen for for ECE in Table 5. For estimating the confidence of Llama’s responses on TriviaQA, our model is better than the best baseline by $11$ percentage points. We are also able to estimate the confidence on the SQuAD dataset using Mistral by $14$ percentage points better than the closest competitor. Qualitatively similar results are seen for the SQuAD dataset using Flan-ul2 (better by $10$ percentage points) and for the CoQA and NQ datasets using Llama (better by $15$ and $10$ percentage points respectively). Our results on the summarization datasets using LLMs that excel at summarization (viz. Pegasus and BART) we see again that we are either better or at least competitive.
@@
Dataset Source LLM Self Target LLM 1 Target LLM 2 Target LLM 3
TriviaQA Llama 0.83 0.74 (Flan-ul2) 0.64 (Mistral) 0.85 (GPT-4)
Flan-ul2 0.74 0.83 (Llama) 0.64 (Mistral) 0.83 (GPT-4)
Mistral 0.64 0.83 (Llama) 0.73 (Flan-ul2) 0.86 (GPT-4)
GPT-4 0.89 0.81 (Llama) 0.71 (Flan-ul2) 0.63 (Mistral)
SQuAD Llama 0.68 0.62 (Flan-ul2) 0.63 (Mistral) 0.81 (GPT-4)
Flan-ul2 0.96 0.89 (Llama) 0.91 (Mistral) 0.82 (GPT-4)
Mistral 0.96 0.90 (Llama) 0.91 (Flan-ul2) 0.81 (GPT-4)
GPT-4 0.83 0.65 (Llama) 0.93 (Flan-ul2) 0.94 (Mistral)
CoQA Llama 0.71 0.79 (Flan-ul2) 0.61 (Mistral) 0.82 (GPT-4)
Flan-ul2 0.80 0.70 (Llama) 0.61 (Mistral) 0.81 (GPT-4)
Mistral 0.61 0.69 (Llama) 0.79 (Flan-ul2) 0.83 (GPT-4)
GPT-4 0.86 0.69 (Llama) 0.79 (Flan-ul2) 0.6 (Mistral)
NQ Llama 0.45 0.46 (Flan-ul2) 0.42 (Mistral) 0.77 (GPT-4)
Flan-ul2 0.47 0.45 (Llama) 0.42 (Mistral) 0.76 (GPT-4)
Mistral 0.42 0.45 (Llama) 0.46 (Flan-ul2) 0.77 (GPT-4)
GPT-4 0.79 0.43 (Llama) 0.46 (Flan-ul2) 0.41 (Mistral)
CNN Pegasus 0.74 0.34 (BART) - -
BART 0.34 0.74 (Pegasus) - -
XSUM Pegasus 0.27 0.34 (BART) - -
BART 0.35 0.25 (Pegasus) - -
@@ 4.2 Confidence Model Interpretability and Transferability
Transferability: Given the commonality between the important features across LLMs for a dataset we now try to test how well does our logistic confidence model for one LLM perform in estimating confidences of another LLM. As seen in Tables 9 and 10 our confidence models are actually quite transferable as they perform comparably or even sometimes better on the other LLMs than the LLM they were built for. This is particularly true for Mistral where, its confidence model performs better for the other two LLMs than itself even coming close in performance to their own confidence models in many cases.
This suggests that we could apply our approach to one LLM and then use the same confidence model to evaluate responses of other LLMs without having to build individual models for them. It would be interesting to further stress test this hypothesis in the future with more LLMs and datasets. Nonetheless, even in the current setup – of six LLMs and six datasets – this observation is interesting and useful.
@@ 5 Discussion
In summary, we have provided an extensible framework for black-box confidence estimation of LLM responses by proposing novel features that are indicative of response correctness. By building an interpretable logistic regression model based on these features we were able to obtain state-of-the-art performance in estimating confidence on six benchmark datasets (CoQA, SQuAD, NQ, TriviaQA, CNN Daily and XSUM) and using six powerful LLMs (Llama-2-13b-chat, Mistral-7B-Instruct-v0.2, Flan-ul2, Pegasus-large and BART-large). The interpretability of our confidence model aided in identifying features (viz. SD, SP, EFA,PP) that were instrumental in driving its performance for different LLM-dataset combinations. This led to the interesting realization that many of the features crucial for performance were shared across the confidence models of different LLMs for a dataset. We thus tested if the confidence models generalized across LLMs for a dataset and found that it indeed was the case leading to the interesting possibility of having an universal confidence model trained on just a single LLMs responses, but applied across many others.
@@
Owing to the supervised nature of training the confidence model, one limitation of our approach is that at least some of the model’s generations must be close to the ground truth for us to obtain a reasonable confidence estimator. Another limitation is that the results and insights were obtained based on datasets in English, but these insights might vary when looking at datasets in other languages. More varied tasks and models could be tested upon in the future. We used rouge to test accuracy of generations consistent with previous works, however, rouge, like also other NLP metrics, can be error prone. Nonetheless, for summarization we also report results with GPT-4 as-a-judge in Tables 13 14 and 15 in the appendix, which are qualitatively similar to those reported in the main paper. In terms of broader impact, our approach can be widely applied as it is simple and works with just black-box access to the LLM. Access to logits or internals of the model are not required. However, our estimates although accurate can be imperfect and this should be taken into account when using our approach in high stakes applications involving LLMs. One should also be cognizant of adversaries aware of our features trying to induce misplaced trust in LLMs they create or prefer.
@@
Given the extensibility of our framework, in the future, it would be interesting to add more features as LLMs evolve. One class of such features might be those where the correctness of a response is checked through creating questions that are (causally) related to the original question and context, and seeing how the response varies by asking this question by itself as opposed to in conjunction with the original question and response. Such and other strategies may help in generalizing these confidence estimators also across datasets something that has been seen when we have additional access to logits of LLMs. Moreover, ideas from selective classification [1, 6] could also be adapted for learning a better confidence model.
@@ Appendix A Prompt Design
  • Flan-ul2 model and GPT-4: Answer the following question in less than 5 words
@@
  • Flan-ul2 model, Llama-2-13b-chat model and GPT-4: Provide an answer in less than 5 words for the following question based on the context below: context: {context} Question: {question} Answer:
@@
  • Flan-ul2 model, Llama-2-13b-chat model and GPT-4: Provide an answer for the following question based on the context below, in less than 5 words:
@@
Prompts for NQ: For all the models we used the following prompt:
@@
Prompt for GPT-4 as-a-judge: Please provide a score between 0 and 1 of how similar the summaries are. 1 indicating very similar and 0 indicating very different.
Table 11: AUROCs on four Q&A and two summarization datasets (CNN, XSUM) using a total of six LLMs (Llama, Flan-ul2, Mistral, GPT-4, Pegasus, BART), where the number of queries to the LLMs is the same for the baselines and our method. Higher values are better. Best results bolded.
Dataset(LLM) # of Lexical EigenValue Eccentricity Degree SE AVC Ours
SS Similarity
TriviaQA(Llama) 0.74 0.76 0.76 0.77 0.77 0.76 0.79 0.88
TriviaQA(Flan-ul2) 0.82 0.81 0.87 0.86 0.86 0.85 0.81 0.95
TriviaQA(Mistral) 0.65 0.72 0.76 0.75 0.75 0.68 0.73 0.81
TriviaQA(GPT-4) 0.89 0.91 0.91 0.92 0.91 0.92 0.94 0.96
SQuAD(Llama) 0.65 0.72 0.74 0.58 0.72 0.61 0.61 0.83
SQuAD(Flan-ul2) 0.6 0.7 0.67 0.65 0.67 0.63 0.66 0.8
SQuAD(Mistral) 0.59 0.7 0.67 0.65 0.67 0.62 0.64 0.84
SQuAD(GPT-4) 0.79 0.82 0.84 0.79 0.83 0.81 0.86 0.91
CoQA(Llama) 0.61 0.74 0.76 0.76 0.77 0.64 0.78 0.92
CoQA(Flan-ul2) 0.61 0.76 0.78 0.78 0.79 0.63 0.76 0.87
CoQA(Mistral) 0.56 0.74 0.79 0.77 0.79 0.59 0.75 0.81
CoQA(GPT-4) 0.81 0.86 0.88 0.87 0.88 0.89 0.91 0.95
NQ(Llama) 0.65 0.75 0.75 0.73 0.74 0.68 0.74 0.85
NQ(Flan-ul2) 0.76 0.76 0.86 0.86 0.86 0.81 0.84 0.93
NQ(Mistral) 0.66 0.73 0.77 0.77 0.78 0.68 0.75 0.83
NQ(GPT-4) 0.81 0.85 0.85 0.85 0.88 0.89 0.9 0.93
CNN (Pegasus) 0.51 0.67 0.73 0.72 0.72 0.55 0.73 0.77
CNN (BART) 0.51 0.59 0.52 0.48 0.54 0.53 0.5 0.57
XSUM (Pegasus) 0.51 0.58 0.69 0.70 0.71 0.54 0.71 0.73
XSUM (BART) 0.51 0.59 0.54 0.52 0.52 0.52 0.53 0.57
Table 12: AUARCs on four Q&A and two summarization datasets (CNN, XSUM) using a total of six LLMs (Llama, Flan-ul2, Mistral, Pegasus, BART), where the number of queries to the LLMs is the same for the baselines and our method. Higher values are better. Best results bolded.
Dataset(LLM) # of Lexical EigenValue Eccentricity Degree SE AVC Ours
SS Similarity
TriviaQA(Llama) 0.76 0.8 0.81 0.8 0.8 0.79 0.8 0.83
TriviaQA(Flan-ul2) 0.7 0.72 0.73 0.73 0.73 0.71 0.72 0.74
TriviaQA(Mistral) 0.55 0.63 0.64 0.64 0.64 0.58 0.63 0.64
TriviaQA(GPT-4) 0.8 0.84 0.84 0.84 0.82 0.84 0.85 0.89
SQuAD(Llama) 0.3 0.36 0.37 0.28 0.36 0.36 0.31 0.68
SQuAD(Flan-ul2) 0.73 0.95 0.83 0.82 0.83 0.78 0.83 0.96
SQuAD(Mistral) 0.72 0.93 0.82 0.82 0.82 0.76 0.83 0.96
SQuAD(GPT-4) 0.7 0.72 0.72 0.63 0.66 0.69 0.71 0.83
CoQA(Llama) 0.56 0.67 0.67 0.67 0.67 0.61 0.66 0.71
CoQA(Flan-ul2) 0.7 0.79 0.8 0.79 0.79 0.73 0.77 0.8
CoQA(Mistral) 0.46 0.62 0.64 0.63 0.64 0.51 0.62 0.61
CoQA(GPT-4) 0.68 0.73 0.72 0.73 0.74 0.72 0.76 0.86
NQ(Llama) 0.37 0.41 0.42 0.41 0.41 0.39 0.42 0.45
NQ(Flan-ul2) 0.41 0.44 0.47 0.46 0.45 0.44 0.45 0.47
NQ(Mistral) 0.32 0.38 0.40 0.40 0.39 0.36 0.39 0.42
NQ(GPT-4) 0.69 0.73 0.74 0.74 0.74 0.73 0.72 0.79
CNN (Pegasus) 0.45 0.51 0.53 0.43 0.52 0.48 0.47 0.74
CNN (BART) 0.21 0.22 0.21 0.21 0.21 0.23 0.23 0.34
XSUM (Pegasus) 0.16 0.17 0.19 0.17 0.17 0.21 0.19 0.27
XSUM (BART) 0.21 0.22 0.20 0.21 0.22 0.23 0.22 0.35
Table 13: AUROCs on two summarization datasets (CNN, XSUM) with GPT-4 as a judge. Higher values are better. Best results bolded.
Dataset(LLM) # of Lexical EigenValue Eccentricity Degree SE AVC Ours
SS Similarity
CNN (Pegasus) 0.54 0.65 0.76 0.77 0.75 0.61 0.75 0.81
CNN (BART) 0.55 0.64 0.55 0.52 0.58 0.56 0.54 0.64
XSUM (Pegasus) 0.56 0.62 0.72 0.74 0.73 0.6 0.75 0.79
XSUM (BART) 0.55 0.63 0.56 0.54 0.55 0.56 0.59 0.61
Table 14: AUARCs two summarization datasets (CNN, XSUM) with GPT-4 as a judge. Higher values are better. Best results bolded.
Dataset(LLM) # of Lexical EigenValue Eccentricity Degree SE AVC Ours
SS Similarity
CNN (Pegasus) 0.49 0.55 0.58 0.49 0.57 0.52 0.53 0.77
CNN (BART) 0.25 0.26 0.27 0.26 0.26 0.27 0.29 0.35
XSUM (Pegasus) 0.19 0.22 0.23 0.2 0.21 0.23 0.21 0.29
XSUM (BART) 0.26 0.26 0.25 0.27 0.27 0.27 0.26 0.37
Table 15: ECEs two summarization datasets (CNN, XSUM) with GPT-4 as a judge. Lower values are better. Best results bolded.
Dataset(LLM) # of Lexical EigenValue Eccentricity Degree SE AVC Ours
SS Similarity
CNN (Pegasus) 0.18 0.14 0.11 0.1 0.09 0.15 0.07 0.05
CNN (BART) 0.48 0.17 0.24 0.25 0.22 0.22 0.22 0.14
XSUM (Pegasus) 0.18 0.18 0.13 0.11 0.09 0.17 0.1 0.06
XSUM (BART) 0.23 0.19 0.21 0.23 0.23 0.22 0.2 0.16